# Phase-shifted remote photoplethysmography for estimating heart rate and blood pressure from facial video

### v3

@@ Abstract
Human health can be critically affected by cardiovascular diseases, such as hypertension, arrhythmias, and stroke. Heart rate and blood pressure are important physiological information for monitoring of cardiovascular system and early diagnosis of cardiovascular diseases. Previous methods for estimating heart rate are mainly based on electrocardiography and photoplethysmography, which require contacting sensors to skin surfaces. Existing cuff-based methods for measuring blood pressure cause inconvenience and are difficult to be utilized in daily life. To address these limitations, this paper proposes a 2-stage deep learning framework for estimating heart rate and blood pressure from facial video. The proposed algorithm consists of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). DRP-Net infers remote photoplethysmography (rPPG) signals at acral and facial sites, and these phase-shifted rPPG signals are utilized to estimate heart rate. BBP-Net integrates temporal features and analyzes phase discrepancy between the acral and facial rPPG signals to estimate systolic blood pressure (SBP) and diastolic blood pressure (DBP). We augmented facial videos in temporal aspects by utilizing a frame interpolation model to increase bradycardia and tachycardia data. Moreover, we reduced blood pressure error by incorporating a scaled sigmoid function in the BBP-Net. Experiments were conducted on MMSE-HR and V4V datasets to demonstrate the effectiveness of the proposed method. Our method achieved the state-of-the-art performance for estimating heart rate and blood pressure with significant margins compared to previous methods. Code is available at https://github.com/GyutaeHwang/phase_shifted_rPPG.
keywords:
@@ 1 Introduction
The cardiovascular system consists of heart and blood vessels, and it circulates blood throughout the body, delivering oxygen and nutrients to tissues while removing waste substances. Abnormalities in the cardiovascular system can lead to various cardiovascular diseases such as cardiomyopathy, hypertensive heart disease, arrhythmias, with potentially severe implications for overall health. Moreover, cardiovascular diseases associated with blood vessels, such as hypertension and stroke, are contribute to increasing the global mortality rate [Gaidai et al., 2023]. However, these diseases can be detected at an early stage through the monitoring of heart rate (HR) and blood pressure (BP) using healthcare devices such as smartwatches. For example, arrhythmias caused by disorders of the sinoatrial node can be prevented through the measurement of bradycardia, tachyarrhythmia, and irregular heart rates. Additionally, systolic blood pressure (SBP) above 120 mmHg and diastolic blood pressure (DBP) above 90 mmHg may indicate the potential presence of hypertension. Recently, interest in cardiovascular diseases has led to growing attention on research for healthcare services [Etemadi et al., 2023, Aggarwal & Pandey, 2023] and monitoring of physiological information [Alnaggar et al., 2023, Gupta et al., 2023].
@@
Heart rate and blood pressure are main physiological information for the monitoring of cardiovascular diseases [Perret-Guillaume et al., 2009, Fuchs & Whelton, 2020], and it can be measured by utilizing various physiological sensors and medical equipment. Heart rate is commonly measured by utilizing electrocardiography (ECG) and photoplethysmography (PPG) sensors. ECG employs electrodes attached to body to record electrical activities caused by contractions and relaxations of the heart. PPG is a non-invasive method that uses a light source and photodetector attached at skin to measure volumetric variations of blood in microvessels. The heart rate can be measured by computing peak-to-peak intervals of physiological signals in the time domain or by analyzing power spectrum in the frequency domain. On the other hand, blood pressure can be measured based on oscillometric methods, which record the magnitude of oscillations using a blood pressure cuff. A catheter-based method is an invasive approach to measure blood pressure, and it involves direct insertion of a sensor into an artery to measure real-time arterial blood pressure (ABP). SBP and DBP can be computed from peak and valley values of the ABP signals. Pulse transit time (PTT) is temporal delay of blood pulse waves which travel from the heart to an acral site such as a fingertip, and it is known that PTT is closely correlated with blood pressure [Geddes et al., 1981]. Although there have been proposed blood pressure estimation methods by analyzing PTT from ECG or PPG signals [Barvik et al., 2021, Ganti et al., 2020], these approaches have intrinsic limitations of requiring skin contacts.
@@
Recently, camera sensors have been utilized to obtain physiological signals, and this approach is called remote photoplethysmography (rPPG). From a facial video, rPPG technique extracts subtle variations in skin color induced by cardiac pulses. The extracted temporal changes in pixel intensities are then transformed into continuous waveforms analogous to conventional PPG signals. The precision of rPPG techniques can be critically affected by many factors such as light conditions, motion artifacts, and different skin tones. Despite these challenges, rPPG has received attention due to non-contact and non-invasive attributes that facilitate remote monitoring of physiological information.
@@
Deep learning methods have been utilized to extract rPPG signals and to estimate heart rate [Chen & McDuff, 2018, Yu et al., 2022, Ouzar et al., 2023]. Additionally, various datasets have been released for the development and evaluation of deep learning algorithms [Bobbia et al., 2019, Niu et al., 2019, Soleymani et al., 2011, Li et al., 2018, Stricker et al., 2014, Zhang et al., 2016b, Revanur et al., 2021]. During collecting the datasets, PPG signals were usually measured at acral sites such as fingertips, and therefore, there exists temporal discrepancy between PPG signals and the corresponding facial videos. However, most previous methods neglected temporal discrepancy between the two modalities; time domain losses have been utilized to measure the difference between PPG signals at acral sites and rPPG signals at facial regions. For example, Lu et al. [2021] and Yu et al. [2022] employed negative Pearson correlation loss to measure the similarity between PPG signals and estimated rPPG signals. In this paper, we introduce the concept of dual rPPG signals, consisting of facial rPPG and acral rPPG signals, to analyze the temporal discrepancy between facial videos and PPG signals. The acral rPPG signal is guided by a time domain loss function, while the facial rPPG signal is guided by frequency domain features which have temporally global characteristics.
@@
On the other hand, blood pressure has been estimated by utilizing PPG signals, multimodal physiological signals, or facial videos. While PPG-based methods have been widely employed in wearable devices, this approach has limitations of requiring physical contact to obtain physiological signals. Although multimodal sensors which measure PPG and ECG signals can accurately estimate blood pressure by analyzing PTT, it also has the disadvantage of requiring contact with ECG and PPG sensors. Recently, camera-based methods have been proposed to estimate blood pressure in a non-contact manner by utilizing spatiotemporal features in facial videos. However, there is a performance gap between the camera-based and PPG-based methods. The objective of this study is to improve the performance of the camera-based approach by utilizing phase-shifted information in acral and facial rPPG signals.
@@
This paper proposes a 2-stage deep learning pipeline consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). In the first stage, DRP-Net infers acral and facial rPPG signals, and these signals are utilized to estimate heart rate. In the second stage, BBP-Net analyzes phase discrepancy between the acral and facial rPPG signals and integrates temporal features based on a multi-scale fusion (MSF) module to estimate SBP and DBP values. A scaled sigmoid function is employed in the BBP-Net to improve the precision of blood pressure estimation by constraining the estimated values into a predefined range. Experiments were conducted on the MMSE-HR database [Zhang et al., 2016b] and V4V database [Revanur et al., 2021] to demonstrate the effectiveness of the proposed method for estimating heart rate and blood pressure from facial videos. The main contributions of this paper can be summarized as follows.
  • 1.
@@
    We propose a novel 2-stage deep learning framework consisting of DRP-Net and BBP-Net for estimating heart rate and blood pressure.
  • 2.
@@
The rest of this paper is organized as follows. Section 2 presents related work. Section 3 explains the proposed method for estimating phase-shifted rPPG signals, heart rate, and blood pressure. Sections 4 and Section 5 present experimental results and conclusion.
2 Related work
2.1 Previous methods for estimating rPPG signals and heart rate
@@
Recently, deep learning methods have been proposed to extract rPPG signals from facial videos. Convolutional neural networks (CNNs) and transformer architectures have been utilized to analyze spatiotemporal features from image sequences. Face detection and spatial attention modules are optionally utilized to improve the robustness to motion artifacts and external brightness conditions. Chen & McDuff [2018] introduced the convolutional attention network (CAN) that employs a Siamese-structured convolutional neural network. This model takes an image frame and the difference map to its adjacent frame, and the spatial attention module analyzes color variations in skin regions. Nowara et al. [2021] introduced an inverse attention module to estimate corrupted signals affected by motion and illumination changes. They further employed Long Short-Term Memory (LSTM) to enhance temporal robustness in estimating physiological signals. Yu et al. [2019] proposed PhysNet3D, which consists of 3D CNN layers for extracting spatiotemporal features and deconvolution layers for recovering temporal details. The PhysNet3D was trained by utilizing the Pearson correlation coefficient loss between PPG and estimated rPPG signals. Yu et al. [2022] proposed a video transformer consisting of temporal difference convolution (TDC) layers. The TDC layers extract local spatial-temporal features to generate query and key projections, and multi-head self-attention mechanism is utilized to integrate global information. Yu et al. [2023] further proposed a SlowFast Network to improve temporal representations of rPPG signals, and they demonstrated promising accuracy in estimating heart rate on cross-domain datasets.
2.2 Deep learning methods for estimating blood pressure
@@
Recently, vision-based methods for estimating blood pressure have received much attention. Most conventional approaches integrated rPPG methods to extract physiological signals and deep learning models to estimate blood pressure from PPG signals. Wu et al. [2022] proposed FS-Net to estimate SBP and DBP values from three-channel rPPG signals and seven physiological indicators including heart rate and body mass index. Bousefsaf et al. [2022] employed continuous wavelet transform and a pre-trained U-Net model to estimate continuous BP signals from estimated rPPG signals. The rPPG signals were obtained by spatially averaging green channel of skin regions in facial videos. On the other hand, Chen et al. [2023] proposed an end-to-end network for estimating blood pressure from facial videos. They extracted spatiotemporal features from four pre-defined facial regions and regressed SBP and DBP values by utilizing ResNet18 and bidirectional LSTM layers.
3 Proposed method
3.1 Overall training pipeline
This paper proposes a two-stage deep learning framework consisting of DRP-Net and BBP-Net to estimate heart rate and blood pressure from facial videos. The proposed deep learning model extracts acral and facial rPPG signals and analyzes their temporal discrepancy to estimate SBP and DBP. Fig. 1 presents an overview of the training pipeline of the proposed method. In the preprocessing step, facial regions are detected within a video clip to define a region of interest (ROI). The DRP-Net learns spatiotemporal features from the ROI sequence and extracts phase-shifted rPPG signals which consists of acral and facial rPPG signals. The BBP-Net consists of multi-scale feature fusion (MSF) blocks and BP prediction heads, and it infers SBP and DBP values from the phase-shifted rPPG signals. For ground truth generation, ABP signals are utilized to calculate pseudo PPG signals, heart rate, SBP, and DBP.
@@ 3.1 Overall training pipeline
Figure 1: Overview of the training pipeline of the proposed method. Red and blue colors indicate ground truth and predicted physiological information, respectively.
3.2 Preprocessing and ground truth generation
@@
The preprocessing step extracts normalized ROI regions from input images to remove redundant background information. The pretrained MTCNN  [Zhang et al., 2016a] model is utilized to detect facial regions, and an ROI is decided to include facial regions within a short video clip. The ROI regions are cropped and resized into the size of 128$\times$128, and their brightness is normalized into the range between 0 and 1 to reduce the effect of light conditions. The cropped ROI sequence is sampled at 25 frames per second (FPS) and split into the window length of 150 frames which corresponds to 6 seconds following the previous method [Chen et al., 2023].
@@
To generate the ground truth data, the ABP signals are synchronized with the facial videos, and they are sampled at 25 Hz. The ABP signals are split into the window length of 6 seconds, and pseudo PPG signals are generated by conducting standardization, detrending, and bandpass filtering. To generate pseudo PPG signals from ABP signals, the detrending algorithm proposed by Tarvainen et al. [2002] is employed for removing the effect of BP variations. Following the previous work [Nowara et al., 2021], bandpass filtering is conducted on physiological signals with a pre-defined range between 0.5 Hz and 3.0 Hz. Fast Fourier Transform is conducted to compute the power spectral density (PSD) of pseudo PPG signals, and the heart rate is computed by analyzing the frequency corresponding to the maximum amplitude of the PSD. Ground truth values for SBP and DBP are computed by averaging peak and valley values of ABP signals within each window.
3.3 Data augmentation

### v4

@@ Abstract
Human health can be critically affected by cardiovascular diseases, such as hypertension, arrhythmias, and stroke. Heart rate and blood pressure are important physiological information for monitoring of cardiovascular system and early diagnosis of cardiovascular diseases. Previous methods for estimating heart rate are mainly based on electrocardiography and photoplethysmography, which require contacting sensors to skin surfaces. Existing cuff-based methods for measuring blood pressure cause inconvenience and are difficult to be utilized in daily life. To address these limitations, this paper proposes a two-stage deep learning framework for estimating heart rate and blood pressure from facial video. The proposed algorithm consists of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). DRP-Net infers remote photoplethysmography (rPPG) signals at acral and facial sites, and these phase-shifted rPPG signals are utilized to estimate heart rate. BBP-Net integrates temporal features and analyzes phase discrepancy between the acral and facial rPPG signals to estimate systolic blood pressure and diastolic blood pressure. We augmented facial videos in temporal aspects by utilizing a frame interpolation model to increase bradycardia and tachycardia data. Moreover, we reduced blood pressure error by incorporating a scaled sigmoid function in the BBP-Net. Experiments were conducted on MMSE-HR and V4V datasets to demonstrate the effectiveness of the proposed method. Our method achieved the state-of-the-art performance for estimating heart rate and blood pressure with significant margins compared to previous methods. Our code is available at https://github.com/GyutaeHwang/phase_shifted_rPPG.
Index Terms:
@@ I Introduction
The cardiovascular system consists of heart and blood vessels, and it circulates blood throughout the body, delivering oxygen and nutrients to tissues while removing waste substances. Abnormalities in the cardiovascular system can lead to various cardiovascular diseases such as cardiomyopathy, hypertensive heart disease, arrhythmias, with potentially severe implications for overall health. Moreover, cardiovascular diseases associated with blood vessels, such as hypertension and stroke, contribute to increasing the global mortality rate [1]. However, these diseases can be detected at an early stage through the monitoring of heart rate (HR) and blood pressure (BP) using healthcare devices such as smartwatches. For example, arrhythmias caused by disorders of the sinoatrial node can be prevented through the measurement of bradycardia, tachyarrhythmia, and irregular heart rates. Additionally, systolic blood pressure (SBP) above 120 mmHg and diastolic blood pressure (DBP) above 90 mmHg may indicate the potential presence of hypertension. Recently, interest in cardiovascular diseases has led to growing attention on research for healthcare services [2, 3] and monitoring of physiological information [4, 5].
@@
Heart rate and blood pressure are main physiological information for the monitoring of cardiovascular diseases [6, 7], and it can be measured by utilizing various physiological sensors and medical equipment. Heart rate is commonly measured by utilizing electrocardiography (ECG) and photoplethysmography (PPG) sensors. ECG employs electrodes attached to body to record electrical activities caused by contractions and relaxations of the heart. PPG is a non-invasive method that uses a light source and photodetector attached at skin to measure volumetric variations of blood in microvessels. The heart rate can be measured by computing peak-to-peak intervals of physiological signals in the time domain or by analyzing power spectrum in the frequency domain. On the other hand, blood pressure can be measured based on oscillometric methods, which record the magnitude of oscillations using a blood pressure cuff. A catheter-based method is an invasive approach to measure blood pressure, and it involves direct insertion of a sensor into an artery to measure real-time arterial blood pressure (ABP). SBP and DBP can be computed from peak and valley values of the ABP signals. Pulse transit time (PTT) is temporal delay of blood pulse waves which travel from the heart to an acral site such as a fingertip, and it is known that PTT is closely correlated with blood pressure [8]. PTT is measured differently depending on the distance between the heart and acral sites, resulting in a temporal delay in the PPG signals at each region. Although there have been proposed blood pressure estimation methods by analyzing PTT from ECG or PPG signals [9, 10], these approaches have intrinsic limitations of requiring skin contacts.
@@
Recently, camera sensors have been utilized to obtain physiological signals through a contactless method called remote photoplethysmography (rPPG). From a facial video, rPPG technique extracts subtle variations in skin color induced by cardiac pulses. The extracted temporal changes in pixel intensities are then transformed into continuous waveforms analogous to conventional PPG signals. The precision of rPPG techniques can be critically affected by many factors such as light conditions, motion artifacts, and different skin tones. Despite these challenges, rPPG has received attention due to non-contact and non-invasive attributes that facilitate remote monitoring of physiological information.
@@
Deep learning methods have been utilized to extract rPPG signals and to estimate heart rate [11, 12, 13]. Additionally, various datasets have been released for the development and evaluation of deep learning algorithms [14, 15, 16, 17]. During collecting the datasets, PPG signals were usually measured at acral sites such as fingertips, and therefore, there exists temporal discrepancy between PPG signals and the corresponding facial videos. However, most previous methods neglected temporal discrepancy between the two modalities; time domain losses have been utilized to measure the difference between PPG signals at acral sites and rPPG signals at facial regions. For example, Lu et al. [18] and Yu et al. [12] employed negative Pearson correlation loss to measure the similarity between PPG signals and estimated rPPG signals. In [19], the authors constructed a multi-site physiological monitoring (MSPM) dataset consisting of full-body PPG signals, verifying the differences in PTT across various body sites. Moreover, Dong et al. [20] found significant trends and phase differences between facial and acral PPG by constructing a contact PPG acquisition system. In this paper, we introduce the concept of phase-shifted rPPG signals, consisting of facial rPPG and acral rPPG signals, to analyze the temporal discrepancy between facial videos and PPG signals. The acral rPPG signal is guided by a time domain loss function, while the facial rPPG signal is guided by frequency domain features which have temporally global characteristics.
@@
On the other hand, blood pressure has been estimated by utilizing PPG signals, multimodal physiological signals, or facial videos. While PPG-based methods have been widely employed in wearable devices, this approach has limitations of requiring physical contact to obtain physiological signals. Although multimodal sensors which measure PPG and ECG signals can accurately estimate blood pressure by analyzing PTT, it also has the disadvantage of requiring contact with ECG and PPG sensors. Recently, camera-based methods have been proposed to estimate blood pressure in a non-contact manner by utilizing spatiotemporal features in facial videos. However, there is a performance gap between the camera-based and PPG-based methods. [21, 22] leverages PTT from different facial regions to extract features related to blood pressure, yet there is a lack of experimental results on publicly available datasets. The objective of this study is to improve the performance of the camera-based approach by utilizing phase-shifted information in acral and facial rPPG signals.
@@
This paper proposes a two-stage deep learning pipeline consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). In the first stage, DRP-Net infers acral and facial rPPG signals, and these signals are utilized to estimate heart rate. In the second stage, BBP-Net analyzes phase discrepancy between the acral and facial rPPG signals and integrates temporal features based on a multi-scale fusion (MSF) module to estimate SBP and DBP values. A scaled sigmoid function is employed in the BBP-Net to improve the precision of blood pressure estimation by constraining the estimated values into a predefined range. Experiments were conducted on the MMSE-HR database [16] and V4V database [17] to demonstrate the effectiveness of the proposed method for estimating heart rate and blood pressure from facial videos. The main contributions of this paper can be summarized as follows.
@@
  • We propose a novel two-stage deep learning framework consisting of DRP-Net and BBP-Net for estimating heart rate and blood pressure.
@@
The rest of this paper is organized as follows. Section II presents related work. Section III explains the proposed method for estimating phase-shifted rPPG signals, heart rate, and blood pressure. Section IV and Section VI present experimental results and conclusion.
II Related work
II-A Estimation of rPPG signals and heart rate
@@
Recently, deep learning methods have been proposed to extract rPPG signals from facial videos. Convolutional neural networks (CNNs) and transformer architectures have been utilized to analyze spatiotemporal features from image sequences. Face detection and spatial attention modules are optionally utilized to improve the robustness to motion artifacts and external brightness conditions. Chen & McDuff [11] introduced the convolutional attention network (CAN) that employs a Siamese-structured convolutional neural network. This model takes an image frame and the difference map to its adjacent frame, and the spatial attention module analyzes color variations in skin regions. Nowara et al. [27] introduced an inverse attention module to estimate corrupted signals affected by motion and illumination changes. They further employed Long Short-Term Memory (LSTM) to enhance temporal robustness in estimating physiological signals. Yu et al. [28] proposed PhysNet3D, which consists of 3D CNN layers for extracting spatiotemporal features and deconvolution layers for recovering temporal details. The PhysNet3D was trained by utilizing the Pearson correlation coefficient loss between PPG and estimated rPPG signals. However, learning rPPG requires a rich temporal representation, which can be a weakness for CNN-based models with limited long-term dependency. To address this issue, Yu et al. [12] proposed a video transformer consisting of temporal difference convolution (TDC) layers. The TDC layers extract local spatial-temporal features to generate query and key projections, and multi-head self-attention mechanism is utilized to integrate global information. Yu et al. [29] further proposed a SlowFast Network to improve temporal representations of rPPG signals, and they demonstrated promising accuracy in estimating heart rate on cross-domain datasets. Although transformer-based models enhance temporal representation with global features, they require high computational complexity.
II-B Deep learning methods for estimating blood pressure
@@
Recently, vision-based methods for estimating blood pressure have received much attention. Most conventional approaches integrated rPPG methods to extract physiological signals and deep learning models to estimate blood pressure from PPG signals. Wu et al. [34] proposed FS-Net to estimate SBP and DBP values from three-channel rPPG signals and seven physiological indicators including heart rate and body mass index. Bousefsaf et al. [35] employed continuous wavelet transform and a pre-trained U-Net model to estimate continuous BP signals from estimated rPPG signals. The rPPG signals were obtained by spatially averaging green channel of skin regions in facial videos. On the other hand, Chen et al. [36] proposed an end-to-end network for estimating blood pressure from facial videos. They extracted spatiotemporal features from four pre-defined facial regions and regressed SBP and DBP values by utilizing ResNet18 and bidirectional LSTM layers. Previous studies typically extract visible pulse waves in a non-parametric manner and have difficulty fully utilizing the temporal features of facial videos. To address this issue, we explore a deep learning-based phase-shifted rPPG estimation method as an intermediate step in BP estimation. In Table I, we summarize previous studies on deep learning-based heart rate and blood pressure estimation.
TABLE I: Summary of Deep Learning Models for HR and BP Estimation.
Methods Tasks Model architecture Input signal
DeepPhys [11] HR 2D CNN Facial video
Benefit of distraction [27] HR 2D CNN and bi-LSTM Facial video
PhysNet [28] HR 3D CNN Facial video
PhysFormer [12] HR Video transformer Facial video
PhysFormer++ [29] HR Video transformer Facial video
Miao et al. [30] BP ResNet and LSTM ECG signal
PP-Net [31] BP, HR 1D CNN and LSTM PPG signal
MLP-BP [32] BP MLP-Mixer PPG and ECG signals
SPT [33] BP Transformer PPG signal
FS-Net [34] BP 2D CNN and FC layers
rPPG signal and
7 physiological indicators
Bousefsaf et al. [35] BP 2D CNN iPPG signal
BPE-Net [36] BP 2D CNN and bi-LSTM Facial video
III Methodology
III-A Overall training pipeline
@@
Figure 1: Overview of the training pipeline of the proposed method. Red and blue texts indicate ground truth and predicted physiological information, respectively. The green, red, and blue arrows represent the preprocessing of input videos, the generation of ground truth, and the post-processing of model outputs, respectively.
This paper proposes a two-stage deep learning framework consisting of DRP-Net and BBP-Net to estimate heart rate and blood pressure from facial videos. The proposed deep learning model extracts acral and facial rPPG signals and analyzes their temporal discrepancy to estimate SBP and DBP. Fig. 1 presents an overview of the training pipeline of the proposed method. In the preprocessing step, facial regions are detected within a video clip to define a region of interest (ROI). The DRP-Net learns spatiotemporal features from the ROI sequence and extracts phase-shifted rPPG signals, which consist of acral and facial rPPG signals. The BBP-Net consists of MSF blocks and BP prediction heads, and it infers SBP and DBP values from the phase-shifted rPPG signals. For ground truth generation, ABP signals are utilized to calculate pseudo PPG signals, heart rate, SBP, and DBP.
Figure 2: Data augmentation process of bradycardia and tachycardia samples.
III-B Preprocessing and ground truth generation
@@
The preprocessing step extracts normalized ROI regions from input images to remove redundant background information. The pretrained MTCNN [37] model is utilized to detect facial regions, and an ROI is decided to include facial regions within a short video clip. Following DeepPhys [11], we use a fixed bounding box for each video with the scaling factor of 1.6 to address missed detections and handle subject movements. The ROI regions are cropped and resized into the size of 128$\times$128, and their brightness is normalized into the range between 0 and 1 to reduce the effect of light conditions. The cropped ROI sequence is sampled at 25 frames per second (FPS) and split into the window length of 150 frames which corresponds to 6 seconds following the previous method [36].
@@
To generate the ground truth data, the ABP signals are synchronized with the facial videos, and they are sampled at 25 Hz. The ABP signals are split into the window length of 6 seconds, and pseudo PPG signals are generated by conducting standardization, detrending, and bandpass filtering (BPF). To generate pseudo PPG signals from ABP signals, the detrending algorithm proposed by Tarvainen et al. [38] is employed for removing the effect of BP variations. Following the previous work [27], BPF is conducted on physiological signals with a pre-defined range between 0.5 Hz and 3.0 Hz. Fast Fourier Transform is conducted to compute the power spectral density (PSD) of pseudo PPG signals, and the heart rate is computed by analyzing the frequency corresponding to the maximum amplitude of the PSD. Ground truth values for SBP and DBP are computed by averaging peak and valley values of ABP signals within each window.
III-C Data augmentation